OpenStudy (youngabudabbin):

What is the product in simplest form? State any restrictions on the variable.

OpenStudy (youngabudabbin):

\[\frac{ y ^{2} }{ y-3 }.\frac{ y ^{2}-y-6 }{ y ^{2+1y} }\]

OpenStudy (youngabudabbin):

@sleepyjess pls help me

OpenStudy (youngabudabbin):

@wolf1728 pls help me

OpenStudy (youngabudabbin):

@ivettef365 pls help me

OpenStudy (youngabudabbin):

pls help me anyone

OpenStudy (wolf1728):

sorry, that's a little bit beyond my knowledge

OpenStudy (youngabudabbin):

@wolf1728 Thanks anyway :)

OpenStudy (mjdennis):

Lots of ways to start. I'd consider trying to factor the term y*2−y−6 and see if you can cancel anything.

OpenStudy (nincompoop):

is that supposed to be an exponent? \(\large y^{\color{red}{2+1y}}\)

jhonyy9 (jhonyy9):

yes always then you need working with fractions with unnknowed terms there are restrictions what will be the value of unknowed term for what value the denominator of fraction(s) will result equal zero - bc. you know that in case that the denominator of a fraction is equal zero so in this case the fraction will be undefined hope this will help you making the restrictions for your exercise easy

jhonyy9 (jhonyy9):

@rebeccaxhawaii

OpenStudy (kinged):

\[\frac{ y+2 }{ y ^{y} }\]

OpenStudy (kinged):

y cannot equal 0.

OpenStudy (mathmate):

@youngAbuDabbin Would you please comment on @nincompoop 's question, namely please doublecheck on the following in red: \(\Huge \frac{ y ^{2} }{ y-3 }.\frac{ y ^{2}-y-6 }{ y ^\color{red}{2+1y} }\) Would it be \(\color{green}{(2+1)y}\), \(\color{blue}{2+1}\), \(\color{purple}{2+y}\), \(\color{red}{2+1y}\), or simply \(\color{magenta}{y+2}\) without the exponentiation?

OpenStudy (mjdennis):

@Kinged, when you help someone, please keep in mind the OpenStudy Code of Conduct. Avoid jumping straight to an answer, especially when someone else is helping out already. http://openstudy.com/code-of-conduct Don't devalue the question/answer process! Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

OpenStudy (kinged):

You just think of \[y ^{2+1y}\]as \[y ^{2} \times y ^{y}\]then you factor \[y ^{2}-y-6\]into \[(y+2)(y-3)\] and finally you simplify by cancellation.